[Brahms-l] Shift report 20050518 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed May 18 2005 - 23:50:49 EDT
Shifter leader was Thomas Materna
Shifters were 


Machine had problems (and STAR too), only started at 9:00 pm
runs 14750-14754
BBR HV tripped, got it back on for last run of this shift, still ok at the moment. An time calibrator was switched ON at the beginning while it should have been turned OFF, this has been (easily) fixed.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed May 18 23:51:02 2005

This archive was generated by hypermail 2.1.8 : Wed May 18 2005 - 23:51:14 EDT